# Timer: Transformers for Time Series Analysis at Scale

### v1

@@ B.1 Pre-training
Based on the constructed UTSD datasets of different sizes and difficulties in the unified single series sequence (S3) format, Timer is pre-trained with increasing data sizes and model parameters to validate the scalability. The detailed model configurations and parameter count of the pre-trained models involved in this paper are shown in Table 4.
Table 4: Detailed model configurations of Timer and corresponding parameter counts. The number of heads for models of all is fixed as $8$.
@@
All experiments were implemented in PyTorch (Paszke et al., 2019) and trained using NVIDIA A100 Tensor Core GPU. We use AdamW (Kingma & Ba, 2015) as the optimizer and cosine annealing algorithm for learning rate decay. The base learning rate is $5\times 10^{-5}$, and the final learning rate is $2\times 10^{-6}$ The decay steps are proportional to the number of training steps of 10 epochs. During pre-training, we use $N=15$ as the number of tokens, and the batch size is set to 8192.
@@ B.2 Downstream Tasks
We introduce the details of downstream experiments and present the generative scheme for each task, including time series forecast, imputation, and anomaly detection. Configurations for downstream adaptation are listed in Table 6. Corresponding detailed results are provided in Section C. And showcases of downstream tasks are shown in Figure 1516, and 17.
@@
The downstream forecasting task is tested on five real-world datasets, including (1) ETT (Zhou et al., 2021) contains 7 variates of power transformers, with the period from July 2016 to July 2018, including four subsets and sampling intervals of one hour and fifteen minutes. (2) ECL (Wu et al., 2021) mainly consists of hourly electricity consumption data from 321 customers (3) Traffic (Wu et al., 2021) collected hourly road occupancy rates measured by 862 sensors on the San Francisco Bay Area highway from January 2015 to December 2016. (4) Weather (Wu et al., 2021) consists of 21 meteorological variates collected every 10 minutes from the Max Planck Institute of Biogeochemistry meteorological station in 2020. (5) PEMS contains California public transportation network data collected through a 5-minute window with the same four common subsets (PEMS03, PEMS04, PEMS07, PEMS08) used in SCINet (Liu et al., 2022).
@@
PEMS04 307 (10172,3375,281) 5min Transportation
PEMS07 883 (16911,5622,468) 5min Transportation
PEMS08 170 (10690,3548,265) 5min Transportation
We adopt the autoregressive generation training objective (Bengio et al., 2000) for downstream forecasting datasets in the fine-tuning stage. Specifically, we divide the lookback length into $N=7$ tokens with the segment length $S=96$. The model naturally outputs $N$ next tokens, which we calculate the mean squared error (MSE) of the $N$ tokens with corresponding ground truth and backpropagate the loss. During inference, we conduct autoregressive generation by concatenating the forecasted result with the lookback series and repeatedly adopting the model to generate the next token until the total length of predicted tokens reaches the forecasted length (if exceeding the predicted length, we crop the result).
@@
$L_{min}$ $L_{max}$ $d_{min}$$^{\dagger}$ $d_{max}$$^{\dagger}$ ${}^{\ast}$ Loss Batch Size Epochs
@@
Considering the prevalent scenarios of anomaly detection by monitoring real-time measurements, the quick judgment of on-the-fly time series anomaly can be more practical in the real world. Therefore, we propose a predictive protocol of anomaly detection based on generative models. Concretely, we use the observed segments to predict the future segment, and the predicted segment will be established as the standard to be compared with the actual value received. We adopt the UCR Anomaly Archive proposed by Wu & Keogh (2021). The task is to find the position of an anomaly in the test series based on a single normal series for training, which is an extremely data-scarce scenario with only one available sample. For downstream adaption, we adopt the same next token prediction as the pre-training, that is, training Timer with the lookback series containing $N=7$ segments of the length $S=96$ to generate the next token with length $96$, which is regarded as the standard value. After training, we record the MSE of all segments in the test set and sort them in descending order. We find the first segment hit the anomaly interval labeled in the dataset within the first $\alpha$ quantile, and we record the quantile. Based on the above protocol, the real-time judgment ability of the model for sudden anomalies can be predictively examined. With more complex time series anomalies introduced by the UCR Anomaly Archive, we hope to establish a more reasonable and challenging benchmark in the field of anomaly detection.
@@ C.1 Time Series Forecasting
Table 7: Full forecasting results of Timer obtained by training from scratch (None) and fine-tuning from UTSD-12G pre-trained model. The bold values we use indicate that the pre-trained model results have positive benefits compared to from scratch. We attach the current state-of-the-art results as SOTA in this table, including PatchTST (Nie et al., 2022) on ETTh1 and Weather, as well as iTransformer (Liu et al., 2023) on ECL, Traffic, PEMS03, and PEMS04. We adopt the unified lookback length as $672$ and the forecast length as $96$.
@@ C.2 Imputation
In this section, we provide the detailed results of the imputation task, including Timer trained from scratch and adapting pre-trained models with $5\%$ available samples in Table 8, $20\%$ samples in Table 9, and full samples in Table 10 on the downstream task. We also report the results of TimesNet at the above three ratios in Table 11. Based on the result, we provided an improvement in imputation performance before and after pre-training with $\{5\%,20\%,100\%\}$ samples in Figure 6 and 14, reflecting the benefits of autoregressive pre-training in segment-wise imputation task.
Figure 14: Pre-training benefit of Timer on downstream imputation task with $20\%$ and $100\%$ available samples.The complete data details used to calculate the pre-training benefits relative to training from scratch in this figure are listed in Table 9 and 10.
@@ C.3 Anomaly Detection
In this section, we provide detailed results of anomaly detection in Table 12, including the results of Timer from scratch and from pre-trained. We conducted experiments on all 250 datasets of UCR Anomaly Archive and calculated the corresponding $\alpha$ quantiles. The results show that the pre-trained Timer can detect time series anomalies with lower $\alpha$ on most datasets.
@@ C.4 Scalability
Table 13: Detailed results for scaling up Timer and the pre-trained dataset.
@@
Table 14: Additional downstream forecasting results of the subsets of PEMS and ETT under different data scarcity of the encoder-only and decoder-only Transformer. The bold part indicates that the result performs best in the current dataset and sample ratio.
@@ Appendix D Showcase
To present a clear performance of our proposed Timer, we provide visualizations for downstream forecasting, imputation, and anomaly detection tasks in Figure 1516 and 17. The forecasting and imputation contain experimental results at different sample ratios, while in anomaly detection, we provide the position of the anomaly and the generated results of our model.

### v2

@@ B.1 Pre-training
Based on the constructed UTSD datasets of different sizes and difficulties in the unified single series sequence (S3) format, Timer is pre-trained with increasing data sizes and model parameters to validate the scalability. Detailed configurations and parameter counts of the pre-trained models involved in this paper are provided in Table 6.
Table 6: Detailed model configurations of Timer and corresponding parameter counts. The number of heads for models is fixed as $8$.
@@
All experiments are implemented in PyTorch (Paszke et al., 2019) and trained using NVIDIA A100 Tensor Core GPU. We use AdamW (Kingma & Ba, 2015) as the optimizer and cosine annealing algorithm for learning rate decay. The base learning rate is $5\times 10^{-5}$, and the final learning rate is $2\times 10^{-6}$ The decay steps are proportional to the number of training steps of 10 epochs. During pre-training, we use $N=15$ as the number of tokens, and the batch size is set to 8192.
@@ B.2 Downstream Tasks
We introduce the details of downstream experiments and present the generative scheme for each task, including time series forecast, imputation, and anomaly detection. Configurations for downstream adaptation are listed in Table 8. Corresponding detailed results are provided in Section C. And showcases of downstream tasks are shown in Figure 1920, and 21.
@@
The downstream forecasting task is tested on the real-world datasets, including (1) ETT (Zhou et al., 2021) contains 7 variates of power transformers, with the period from July 2016 to July 2018, including four subsets and sampling intervals of one hour and fifteen minutes. (2) ECL (Wu et al., 2021) mainly consists of hourly electricity consumption data from 321 customers (3) Traffic (Wu et al., 2021) collected hourly road occupancy rates measured by 862 sensors on the San Francisco Bay Area highway from January 2015 to December 2016. (4) Weather (Wu et al., 2021) consists of 21 meteorological variates collected every 10 minutes from the Max Planck Institute of Biogeochemistry meteorological station in 2020. (5) PEMS contains California public transportation network data collected through a 5-minute window with the same four common subsets (PEMS03, PEMS04, PEMS07, PEMS08) used in SCINet (Liu et al., 2022).
@@
      PEMS04       307       (10172,3375, 3375)       5min       Transportation
      PEMS07       883       (16911, 5622, 5622)       5min       Transportation
      PEMS08       170       (10690, 3548, 3548)       5min       Transportation
We adopt the autoregressive generation training objective (Bengio et al., 2000) for downstream forecasting datasets in the fine-tuning stage. Specifically, we divide the lookback length into $N=7$ tokens with the segment length $S=96$. The model naturally outputs $N$ next tokens, which we calculate the mean squared error (MSE) of the $N$ tokens with corresponding ground truth and backpropagate the loss. During inference, we conduct iterative multi-step forecasting by concatenating the forecasted result with the lookback series and repeatedly adopting the model to generate the next token until the total length of predicted tokens reaches the expected length. If exceeding the predicted length, we will crop the excess value of the end.
@@
$L_{min}$ $L_{max}$ $d_{min}$ $d_{max}$ LR Loss Batch Size Epochs
@@
Considering the prevalent scenarios of anomaly detection by monitoring real-time measurements, the quick judgment of on-the-fly time series anomaly can be more practical in the real world. Therefore, we propose a predictive protocol of anomaly detection based on generative models. Concretely, we use the observed segments to predict the future segment, and the predicted segment will be established as the standard to be compared with the actual value received. We adopt the UCR Anomaly Archive proposed by Wu & Keogh (2021). The task is to find the position of an anomaly in the test series based on a single normal series for training, which is an extremely data-scarce scenario with only one available sample. For downstream adaption, we adopt the same next token prediction as the pre-training, that is, training Timer with the lookback series containing $N=7$ segments of the length $S=96$ to generate the next token with length $96$, which is regarded as the standard value. After training, we record the MSE of all segments in the test set and sort them in descending order. We find the first segment hit the anomaly interval labeled in the dataset within the first $\alpha$ quantile, and we record the quantile. Based on the above protocol, the real-time judgment ability of the model for sudden anomalies can be predictively examined. Detailed quantiles of Timer in $250$ tasks are provided in Table 15. With more complex time series anomalies introduced in UCR Anomaly Archive, we hope to establish a reasonable and challenging benchmark in the field of anomaly detection.
Zero-shot forecasting
We conduct zero-shot forecasting experiments on seven datasets from iTransformer (Liu et al., 2023b). Notably, PEMS datasets are not included, as they have already appeared in the LOSTA dataset for pre-training. We apply the same data-split strategy as Autoformer (Wu et al., 2021) and calculate the averaged MSE of all predict-$96$ windows in the test split. We evaluate five open-source large time series models, including Timer, Moiria (Woo et al., 2024), TimesFM (Das et al., 2023b), Chronos (Ansari et al., 2024), and MOMENT (Goswami et al., 2024). We further assess the qualities in Table 9, which includes more LTSMs and summarizes several attributes and abilities of large models.
Table 9: Quality evaluation of large time series models. Architecture denotes the Transformer category. Model size presents the parameter counts. Token type presents the graininess of time series tokens. Context length means the maximum/fixed input length of the model.
Method Timer Moirai MOMENT Chronos Lag-LLama TimesFM TimeGPT-1
(Ours) (2024) (2024) (2024) (2023) (2023b) (2023)
Architecture Decoder Encoder Encoder Encoder Decoder Decoder Encoder
Decoder Decoder Decoder
Model size 29M, 50M, 14M, 91M, 40M, 125M 20M, 46M, 200M 17M, 70M, UnKnown
67M 311M 385M 200M, 710M 200M
Supported tasks Forecast Forecast
Forecast Imputation
Forecast Forecast Forecast Forecast
Imputation
Classification
Detection
Detection
Detection
Pre-training Scale 28B 27.65B 1.13B 84B 0.36B 100B 100B
Token type Segment Segment Segment Point Point Segment Segment
Context length $\leq$1440 $\leq$5000 = 512 $\leq$512 $\leq$1024 $\leq$512 Unknown
Variable length True True False True True True True
Probabilistic False True False True True True True
@@ C.1 Time Series Forecasting
Table 10: Full forecasting results of Timer obtained by training from scratch (None) and fine-tuning from UTSD-12G pre-trained model. The bold values we use indicate that the pre-trained model results have positive benefits compared to from scratch. We attach the current state-of-the-art results as SOTA in this table, including PatchTST (Nie et al., 2022) on ETTh1 and Weather, as well as iTransformer (Liu et al., 2023b) on ECL, Traffic, PEMS03, and PEMS04. We adopt the unified lookback length as $672$ and the forecast length as $96$.
@@ C.2 Imputation
In this section, we provide the detailed results of the imputation task, including Timer trained from scratch and adapting pre-trained models with $5\%$ available samples in Table 11, $20\%$ samples in Table 12, and full samples in Table 13 on the downstream task. We also report the results of TimesNet at the above three ratios in Table 14. Based on the result, we provided an improvement in imputation performance before and after pre-training with $\{5\%,20\%,100\%\}$ samples in Figure 8 and 18, reflecting the benefits of autoregressive pre-training in segment-wise imputation task.
Figure 18: Pre-training benefit of Timer on the downstream imputation task with $20\%$ and $100\%$ available samples. Complete results details used to calculate the pre-training benefits relative to training from scratch in the figures are listed in Table 11-14.
@@ C.3 Anomaly Detection
In this section, we provide detailed results of anomaly detection in Table 15, including the results of Timer from scratch and pre-trained. We conducted experiments on all $250$ datasets of UCR Anomaly Archive and calculated the corresponding $\alpha$ quantiles. The results show that the pre-trained Timer can detect time series anomalies with smaller $\alpha$ on most datasets.
@@ C.4 Scalability
Table 16: Detailed results for scaling up the pre-trained scale and the parameter of Timer.
@@
Table 17: Additional downstream $672$-pred-$96$ forecasting results of the subsets of PEMS and ETT under different data scarcity of the encoder-only and decoder-only Transformer. The bold part indicates that the result performs best in the current dataset and sample ratio.
@@
C.5 Zero-shot forecasting
In this section, we provide detailed results of zero-shot forecasting in Table 18. We conducted experiments on seven datasets that are not included in the pre-training corpora of LTSMs. The results show that the top-ranked LTSMs are Timer, Moiria, and TimesFM. The performance of probabilistic forecaster Chronos can be improved by sampling more trajectories. It can still be an issue that scaling behavior is not evident on some datasets in the zero-shot scenario, and the failure of multi-step prediction can also appear in some models, indicating the development of zero-shot LTSMs is still in the early stage.
Table 18: Zero-shot forecasting evaluation. We extensively evaluate available large time series models. We provided the average rank on all downstream datasets, where the lower is better. For the probabilistic forecaster Chronos: S1 means sampling one trajectory and S20 means sampling 20 trajectories and using the average. ’-’ indicates that multi-step error accumulation leads to failure predictions.
Model
Timer-1B
Timer-16B
Timer-28B
Moirai-S
Moirai-M
Moirai-L
MOMENT
TimesFM
Chronos-S1
Chronos-S20
ETTh1 0.438 0.364 0.393 0.441 0.383 0.394 0.674 0.414 0.571 0.454
ETTh2 0.314 0.294 0.308 0.295 0.295 0.293 0.330 0.318 0.423 0.326
ETTm1 0.690 0.766 0.420 0.562 0.448 0.452 0.670 0.354 0.632 0.451
ETTm2 0.213 0.234 0.247 0.218 0.225 0.214 0.257 0.201 0.272 0.190
ECL 0.192 0.139 0.147 0.212 0.162 0.155 0.744 - - -
Traffic 0.458 0.399 0.414 0.616 0.425 0.399 1.293 - - -
Weather 0.181 0.203 0.243 0.195 0.197 0.221 0.255 - - -
Rank (Avg.) 1.571 2.286 4.429 2.250 5.500 3.250
@@ Appendix D Showcase
To present a clear performance of our proposed Timer, we provide visualizations for downstream forecasting, imputation, and anomaly detection tasks in Figure 1920 and 21. The forecasting and imputation contain experimental results at different sample ratios. For anomaly detection, we provide the position of the anomaly and the generated normal series by Timer.
@@
Appendix E Limitations
UTSD is constructed with hierarchical capacities. Though it is helpful to study the scalability of the model, it is not big enough since we have witnessed recent work claims the pre-training on ten and even a hundred billion time points. Therefore, we advocate for the ongoing expansion of data infrastructure while upholding high quality and hierarchy, which may significantly advance the time series community. In terms of the method, this work aims at an early but important development of large models. Despite the generalization, scalability, and task-generality that Timer has achieved, time series classification has not been unified in our generative formulations and Timer does not yet support probabilistic forecasting and specially adapts for multiple variables. It also leaves for better zero-shot generalization and advanced abilities, such as in-context learning and multi-modality, which are scheduled to be developed by ever-large pre-training.
Appendix F Societal Impacts
Real-world applications
This paper develops large models for the field of time series. We present a general-purpose time series analysis model to handle data-scarce scenarios. Given the state-of-the-art performance of Timer, this model may be applied to many real-world applications, which helps our society prevent risks in advance and make better decisions with limited available samples. Our paper mainly focuses on scientific research and has no obvious negative social impact.
Academic research
In this paper, we release a high-quality dataset for scalable pre-training. Different from prior works, the dataset is not merely aggregation but follows deftly curation. Based on it, the research on scalable time series architectures and pre-training techniques can be facilitated. Towards large time series models, the proposed Timer shows its generalization and versatility in many tasks. The regime of generative pre-training and autoregression can be instructive for future research.